OpenStudy (anonymous):

Match each scenario with the experimental probability that it will occur. Monday: 6, 1, 2, 8 Wednesday: 5, 2, 4 Friday: 5, 8, 1 1) more than 2 people at the track 2) less than 5 people at the track 3) no more than 2 people at the track 4) At least 6 people at the track A) 60% B) 40% C) 50% D) 30%

OpenStudy (anonymous):

I believe I finally answered my own question: 1) = 60 % because there are a total of 10 weeks and out of the 10 weeks there was 6 weeks that there was more than 2 people at the track so 6/10 = 60/100 = 60 % 2) = 50% because there are a total of 10 weeks and out of the 10 weeks there was 5 weeks that there was less than 5 people at the track so 5/10 = 50/100 = 50% 3) = 40% same reason, etc for the rest.